sales at the restaurantA were 450,000 more than B. Total sales were 625,000,000. What are the sales at each store
1
Expert's answer
2013-07-15T08:31:25-0400
Let sales at the store B are x, then sales at the store A are x+450,000. So we have
x+x+450,000=625,000,000
whence 2x=625,000,000-450,000=624,550,000
and therefore x=312,275,000 - sales at B
sales at A are x+450,000=312,725,000